Question 18 (5 points)

 

A concave mirror has a focal length of -27.0 cm. Find the magnification produced by the mirror when the object distance is 9.0 cm and 18.0 cm.

Question 18 options:


A) 

No option correct


B) 

0.750 and 0.600


C) 

0.850 and 0.500


D) 

0.950 and 0.800


E) 

0.650 and 0.700

1
Expert's answer
2021-11-15T10:34:01-0500
"\frac 1f=\frac1{i_1}+\frac1{o_1},\\\space\\
\frac 1{-27}=\frac1{i_1}+\frac1{9},\\\space\\
i_1=-6.75,\\\space\\
m_1=\frac{i_1}{o_1}=-0.75."

Thus, neither option is correct.